# Next-Token Prediction Task Assumes Optimal Data Ordering for LLM Training in Proof Generation

### v1

@@ 4.2 Metrics
Metrics for the intuitionistic propositional logic theorem-proving task.  For the theorem-proving task, we evaluate model performance using two key metrics after prompting the model with a theorem statement.The first metric is tactic prediction correctness, which assesses whether the tactics generated by the model can successfully prove the theorem. The second metric is proof generation correctness, which not only checks if the generated tactics prove the theorem, but also ensures that all intermediate states are correctly generated. The entire evaluation process is automated using Lean. We input the theorem statement and the model-generated tactics into Lean. Lean computes a state after applying the given tactic to the given state. We then compare the model-generated intermediate states with the states computed by Lean. Lean automatically determines whether the tactics successfully prove the theorem.
Metrics for the multiplication task.  In contrast to prior works, we treat 4-digit by 4-digit computation as a proof generation task, not as a computation task Liu and Low (2023); Zhou et al. (2023b, 2024); Lee et al. (2023). We evaluate model performance using the proof correctness metric, final answer correctness metric, and Step 1 correctness metric, where Step 1 refers to the first step in the proof according to the optimal order in the multiplication task. For the proof correctness metric, our model is prompted with only the problem, “Tell me what num1*num2 is and prove it”, and a successful generation of the model requires every proof step in the generation to be correct, together with the final answer. It is in parallel with the usual proof generation scenario, where one is given a statement and is asked to generate the proof to prove whether the statement is true or false. In mathematical reasoning, simply giving out a true or false answer is not enough without correct convincing arguments. For the final answer prediction task, we prompt the model in two ways. The first prompt contains only the problem. The second prompt contains both the problem and all the intermediate proof steps except the final answer correspondingly. For both prompts, we check the correctness of the final answer. For Step 1 correctness metric, we prompt the model with the problem and the intermediate proof steps, except for Step 1. We then check the correctness of Step 1 output by the model.
@@ 4.3 Models and hyperparameters
We use Gemma-2B Team et al. (2024) and Llama-2-7B-hf Touvron et al. (2023) as our proof generators and the default pre-trained tokenizers for both of the models, even in train-from-scratch experiments. We use the function AutoModelForCausalLM.from_config from Huggingface to randomly initialize the models. Unless otherwise stated, the models that we fine-tune are pre-trained with default weights. The learning rate of all training processes is set to be $5\times 10^{-5}$. We use Adam as our optimizer Diederik (2014). All training processes are performed on 4 NVIDIA A100, with the per-device batch size set to 1. This makes the batch size to be 4 for all experiments. All experiments are completed within 24 hours, except for the train-from-scratch experiments, which take about 3 days to finish. Every training process is based on prediction with cross-entropy loss.
During the experiments, for the multiplication task, we fine-tune Gemma-2b and Llama-2-7b-hf on SEQ, PSER, and SER datasets, resulting in a total of six models in this case. We also train from scratch randomly initialized Gemma-2b on SEQ, PSER, and SER as well so that we have three additional models. We refer to models that are fine-tuned/trained from scratch on SEQ, PSER, and SER datasets as SEQ models (3 in total), PSER models (3 in total), and SER models (3 in total).
@@
Figure 2: Experimental results on the multiplication task measured by the metrics of proof generation correctness and final answer correctness. SER models achieves 0 percent success rate for all cases. In (a), (b) and (c), models are prompted with the problems only and performances are measured by the proof correctness metric. In (d), (e) and (f), SEQ and PSER models are prompted with both the problems and the intermediate steps, whereas SER model is prompted with the problems only. Performances are measured by the final answer correctness metric. Results (especially those for SER models) show no evidence for the claim that model can look forward during training.
5 Main Results and Analysis
@@
Zero-shot performance of Gemma-2b and Llama-2-7b-hf (no need for model name) on 4-by-4 multiplication and theorem-proving tasks.  We begin by evaluating the zero-shot performance of both Gemma-2b and Llama-2-7b-hf on two tasks. For the theorem-proving task, we prompt Llama-2-7b-hf with a theorem statement and request it to generate a proof in Lean, see Figure 3f. The model, however, fails to produce any correct solution measured by either the tactic prediction metric or the proof correctness metric, achieving a success rate of 0 out of 1000 test cases. For the multiplication task, we conduct tests using 100 cases. The models are provided with two types of prompts. The first prompt is, “Tell me what num1 * num2 is and prove it." The second prompt incorporates a chain-of-thought (CoT) instruction: “Tell me what num1 * num2 is and prove it. Let’s think step by step." Despite these variations, both pre-trained Gemma-2b and Llama-2-7b-hf fail on the 4-by-4 multiplication task with a success rate of 0/1000 based on proof correctness and final answer accuracy. This result aligns with expectations, as the 4-by-4 multiplication task poses significant challenges even for more advanced models like GPT-4 Liu and Low (2023).
@@
Main observation: Higher right-to-left step-wise dependency of data lead to higher training efficiency on proof generation.  We conduct experiments on both the intuitionistic propositional logic theorem-proving task and the multiplication task. For all the nine models that are fine-tuned/trained from scratch for the multiplication task, we prompt the models with “Tell me what is num1*num2 and prove it”, where (num1, num2) is in the testing data. See Figure 3a, 3b, 3c for prompting configuration. We then measure the generated output of the model by the proof correctness metric. For the intuitionistic propositional logic theorem-proving task, we prompt the SEQ and SER models with the theorem statement. See Figure 3f for prompting configuration. We use both the tactic prediction metric and the proof correctness metric to measure model performance for this task. As shown in Figure 2(a)2(b)2(c)5(a)5(b), SEQ models learn the task a lot more efficiently than the models PSER models and SER models for the multiplication task. For the propositional logic theorem-proving task, the SEQ model also outperforms the SER model by 20 percent measured by the tactic prediction metric. Note that for the multiplication task, for all three base models, PSER models do get a small number of test problems right in the second half of the training, and SER models failed all test data consistently. This performance difference is in line with our claim that putting more intermediate supervision on the left will improve model training efficiency in reasoning.
@@
LLMs don’t look forward in training: higher right-to-left step-wise dependency of data lead to higher training efficiency on final answer prediction.  For this experiment, we focus on the multiplication task. We use the same 9 models as in our first experiments, but for SEQ and PSER models, we prompt the model with both the problems and the intermediate steps excluding the final answer, so that we exclude the snow-balling effect. See Figure 3d, 3e for prompting configuration. For SER models, since there is no intermediate step before the final answer in the data, we prompt it with the problem only. See Figure 3c for prompting configuration in this case. We use final answer prediction correctness for this experiment to measure model performance. As displayed in Figure 2(d)2(e)2(f), in this setting, for all three base models, models that are fine-tuned/trained from scratch on SER got 0 final answers correct consistently. The results of SER show that LLMs don’t look forward during training, i.e. fail to learn to use intermediate supervision on the right in training.
@@
We also note that when both the problem and all intermediate proof steps (except for the final answer) are provided as prompts to the models, SEQ models still outperform PSER models in terms of final answer correctness, as shown in Figure 2(d)2(e)2(f). See Figure 3d, 3e for prompting configuration. By giving the correct intermediate steps as prompts to both SEQ and PSER models, we eliminate the snowballing effect during the generation process. This result is a another evidence showing that stronger right-to-left step-wise dependencies in the data enhance model learning.
@@
SER Models learn spurious dependency.  In this experiment, we test our SEQ and SER models for the multiplication task. The testing prompt of the model is the problem only, and the models are expected to generate the full proof. See Figure 3a, 3c. We use the Step 1 correctness metric, i.e. checking the correctness of Step 1 in the models’ generated output. Step 1 is the decomposition step that splits the 4-digit number into simple terms, e.g. 2345 to 2*1000+3*100+4*10+5*1. During the training, the model should be able to learn this step quickly given the problem due to its simplicity. Indeed, from Figure 4(a)4(b), we see that for SEQ models based on pre-trained models, they learn Step 1 almost perfectly after the first 2500 steps. For the randomly initialized Gemma-2b trained from scratch on SEQ, the model performance measured by the Step 1 correctness metric quickly reaches more than 90% and is approaching perfect scores followed by some fluctuations, as shown by Figure 4(c). For SER models, however, they converge much slower measured by the Step 1 correctness metric in the early training stage compared to SEQ models, as shown in Figure 4(a)4(b)4(c). This slow convergence in the early training stage shows that SER models try hard to fit the spurious dependency of Step 1 on all other proof steps at first. They need more training to realize that Step 1 does not depend on all the other proof steps.
@@ 6 Other Results and Discussion
Pre-train data is not the major cause of the order effect.  One might argue that the order bias in pre-training dataset is the primary cause of the order effects observed in this study. However, for the multiplication task, all of our experiments mentioned above were conducted on both pre-trained models and randomly initialized models, thereby ruling out this possibility given the results.
SEQ is better than PSER even when problem and proof steps are masked during training.  In this experiment, we don’t use our 9 models mentioned before. We fine-tune pre-trained Gemma-2b on SEQ and PSER but with all tokens except final answers masked out. Therefore the only input-output pair that enters into the loss function is (problem+intermediate proof steps, final answer). During testing, we prompt both models with the problem and the intermediate proof steps. See Figure 3d,3e for the prompting configuration. Figure 8(b) shows that the model fine-tuned on masked SEQ outperforms the model fine-tuned scratch on masked PSER measured by final answer correctness for the pre-trained case.
@@
Discussion: how to find the optimal order.  In the context of our multiplication task, identifying the optimal order is theoretically straightforward: one could enumerate all possible proof orders, train the model on each, and then compare the training loss or success rate using appropriate metrics. The order that yields the smallest training loss can be considered optimal. This approach works because all training samples in multiplication share a common structure, e.g. Step 1 decompose, Step 2 distribute and so on. However, this is not the case for more general datasets, where each data sample possesses its own unique structure. A potential solution is to train a separate model to reorder the data entries into an optimal sequence. The key challenge here is to identify the intermediate supervision structure for each data sample, i.e. for a given token, which other tokens within the same sample contribute to predicting it.
@@ 7 Conclusion and Future work
In conclusion, our study highlights the significant impact of order for each data sample on LLMs’ training efficiency for the proof generation tasks. We claim that the optimal order of a data sample, which we call the intuitively sequential order, is to put intermediate supervision toward a particular proof step on the left for all proof steps in a data sample. We show the two causes for this order effect: given suboptimal data order, training inefficiency arises from (i) inability to look forward during training, i.e. learning to use future intermediate supervision to predict the current token, (ii) propensity to learn spurious token dependencies in the early training stage. Our experiments on intuitionistic propositional logic theorem-proving and 4-by-4-digit multiplication provide strong support for these explanations, highlighting the importance of finding the optimal data order in model training.
A promising direction for future work is to investigate this claim during the pre-training stage of LLMs for reasoning tasks. We hypothesize that the claim will hold in this context as well, given that both fine-tuning on downstream tasks and pre-training on large corpora are fundamentally based on the next-token prediction task for decoder-only LLMs. Another future work is to employ machine learning methods to estimate and discover optimal data order, specifically, to identify for every token which other tokens in a data sample provide intermediate supervision. This holds significance on its own, independent of the gain it can bring to model training. For example, suppose the true supervision structure is discovered for mathematical proofs and those proofs are reordered accordingly. In that case, it might enable students to learn more quickly and reveal new mathematical intuition.
@@ Appendix A Appendix
A.1 A graduate-level example of a math theorem proof that follows the logic sequential order instead of the optimal order
We use Chapter 10, Proposition 3.6 (2) in a popular graduate textbook called Abstract Algebra, as an easy example (Dummit and Foote (2004)) to show why the purely logical order is often not the optimal order. Note that the proof here is stated in the purely logical order.
An graduate-level example Theorem: Let $Q$ be an $R$-module: If $R$ is a P.I.D, then $Q$ is injective if and only if $rQ=Q$ for every nonzero $r\in R$. In particular, a $\mathbb{Z}$-module is injective if and only if it is divisible. Proof: Suppose $R$ is a P.I.D. Any nonzero ideal $I$ of $R$ is of the form $I=(r)$ for some nonzero element $r$ of $R$. An $R$-module homomorphism $f:I\mapsto Q$ is completely determined by the image $f(r)=q$ in Q. This homomorphism can be extended to a homomorphism $F:R\mapsto Q$ if and only if there is an element $q^{\prime}$ in $Q$ with $F(a)=q^{\prime}$ satisfying $q=f(r)=F(r)=rq^{\prime}$. It follows that Baer’s criterion for $Q$ is satisfied if and only if $rQ=Q$, which proves the first two statements in (2). Baer’s Criterion: The module $Q$ is injective if and only if for every left ideal $I$ of $R$ any $R$-module homomorphism $g:I\mapsto Q$ can be extended to an $R$-module homomorphism $G:R\mapsto Q$. Change to figure, explain everything
A.1.1 Analysis of the example
In order to prove the theorem, the first natural instinct, after reading the statement, would be to apply Baer’s criterion to prove the theorem. This is because Baer’s criterion list one condition for $Q$ to be injective, and that condition seems to be related to a P.I.D ring $R$ well. Assuming this approach, one can easily construct the proof by extending maps on each left ideal $I$ of $R$ to a homomorphism from $R$ to $Q$, as $R$ is a P.I.D, hence finished the proof. However, without recognizing the necessity of Baer’s criterion from the outset, the proof would be significantly more challenging to write up. Applying Baer’s criterion act as the intermediate supervision toward the first few steps of the proof.
A.2 Data Examples in training

### v2

@@ 4.2 Metrics
Metrics for the intuitionistic propositional logic theorem-proving task.  For the theorem-proving task, we evaluate model performance using two key metrics. The first metric is tactic prediction correctness, which assesses whether the tactics generated by the model can prove the theorem. The second metric is proof generation correctness, which not only checks if the generated tactics prove the theorem, but also ensures that all intermediate states are correctly generated. The entire evaluation process is automated using Lean. We input the theorem statement and the model-generated tactics into Lean. Lean computes a state after applying the given tactic to the given state. We then compare the model-generated intermediate states with the states computed by Lean. Lean automatically determines whether the tactics successfully prove the theorem.
Metrics for the multiplication task.  In contrast to prior works, we treat 4-digit by 4-digit computation as a proof generation task, not as a computation task (Liu and Low, 2023; Zhou et al., 2023b, 2024; Lee et al., 2023). We evaluate model performance using the proof correctness metric, final answer correctness metric, and Step 1 correctness metric, where Step 1 refers to the first step in the proof according to the optimal order in the multiplication task. For the proof correctness metric, our model is prompted with only the problem, “Tell me what num1*num2 is and prove it”, and a successful generation of the model requires every proof step in the generation to be correct, together with the final answer. It is in parallel with the usual proof generation scenario, where one is given a statement and is asked to generate the proof to prove whether the statement is true or false. In mathematical reasoning, simply giving out a true or false answer is not enough without correct convincing arguments. For the final answer prediction task, we prompt the model in two ways. The first prompt contains only the problem. The second prompt contains both the problem and all the intermediate proof steps except the final answer correspondingly. For both prompts, we check the correctness of the final answer. For Step 1 correctness metric, we prompt the model with the problem and the intermediate proof steps, except for Step 1. We then check the correctness of Step 1 output by the model.
@@ 4.3 Models and hyperparameters
We use Gemma-2B (Team et al., 2024) and Llama-2-7B-hf (Touvron et al., 2023) as our proof generators and the default pre-trained tokenizers for both of the models, even in train-from-scratch experiments. We use the function AutoModelForCausalLM.from_config from Huggingface to randomly initialize the models. Unless otherwise stated, the models that we fine-tune are pre-trained with default weights. The learning rate of all training processes is set to be $5\times 10^{-5}$. We use Adam as our optimizer (Diederik, 2014). All training processes are performed on 4 NVIDIA A100, with the per-device batch size set to 1. This makes the batch size to be 4 for all experiments. All experiments are completed within 24 hours, except for the train-from-scratch experiments, which take about 3 days to finish. Every training process is based on prediction with cross-entropy loss.
During the experiments, for the multiplication task, we fine-tune Gemma-2b and Llama-2-7b-hf on SEQ, PSER, and SER datasets, resulting in a total of six models. We also train from scratch randomly initialized Gemma-2b on SEQ, PSER, and SER as well so that we have additional three models. We call models that are fine-tuned/trained from scratch on SEQ, PSER, and SER datasets as SEQ models (3 in total), PSER models (3 in total), and SER models (3 in total) respectively.
@@
Figure 2: Experimental results on the multiplication task measured by the metrics of proof generation correctness and final answer correctness. SER models achieves 0 percent success rate for all cases. In (a), (b) and (c), models are prompted with the problems only and performances are measured by the proof correctness metric. In (d), (e) and (f), SEQ and PSER models are prompted with both the problems and the intermediate steps, whereas SER model is prompted with the problems only. Performances are measured by the final answer correctness metric. Results show no evidence for the claim that the model can look forward during training.
@@
Zero-shot performance of the models on the multiplication and theorem-proving tasks.  We begin by evaluating the zero-shot performance of both Gemma-2b and Llama-2-7b-hf on two tasks with 3-shots examples. For the theorem-proving task, we prompt Llama-2-7b-hf with a theorem statement and request it to generate a proof in Lean, see Figure 5f. The model, however, fails to produce any correct solution measured by either the tactic prediction metric or the proof correctness metric, achieving a success rate of 0 out of 1000 test cases. For the multiplication task, we conduct tests using 100 cases. The models are provided with two types of prompts. The first prompt is, “Tell me what num1 * num2 is and prove it." The second prompt incorporates a chain-of-thought (CoT) instruction: “Tell me what num1 * num2 is and prove it. Let’s think step by step." Both two prompts also include 3 examples in their corresponding formats. Despite these variations, both pre-trained Gemma-2b and Llama-2-7b-hf fail on the 4-by-4 multiplication task with a success rate of 0/1000 based on proof correctness and final answer accuracy. This result aligns with expectations, as the 4-by-4 multiplication task poses significant challenges even for more advanced models like GPT-4 (Liu and Low, 2023).
@@
Main observation: training efficiency is maximized when models are trained on intuitive sequential order.  We conduct experiments on both the intuitionistic propositional logic theorem-proving task and the multiplication task. For all the nine models that are fine-tuned/trained from scratch for the multiplication task, we prompt the models with “Tell me what is num1*num2 and prove it”, where (num1, num2) is in the testing data. See Figure 5a, 5b, 5c for prompting configuration. We then measure the generated output of the model by the proof correctness metric. For the intuitionistic propositional logic theorem-proving task, we prompt the SEQ and SER models with the theorem statement. See Figure 5f for prompting configuration. We use both the tactic prediction metric and the proof correctness metric to measure model performance for this task. As shown in Figure 2(a)2(b)2(c)4(a)4(b), SEQ models learn the task a lot more efficiently than the models PSER models and SER models for the multiplication task. For the propositional logic theorem-proving task, the SEQ model also outperforms the SER model by 20 percent measured by the tactic prediction metric. Note that for the multiplication task, for all three base models, PSER models do get a small number of test problems right in the second half of the training, and SER models failed all test data consistently. This performance difference is in line with our claim that putting more intermediate supervision on the left will improve model training efficiency in reasoning.
@@
LLMs don’t look forward in training.  For this experiment, we focus on the multiplication task. We use the same 9 models as in our first experiments, but for SEQ and PSER models, we prompt the model with both the problems and the intermediate steps excluding the final answer, so that we exclude the snow-balling effect (the effect that the earlier errors model makes will affect model performance in later outputs). See Figure 5d, 5e for prompting configuration. For SER models, since there is no intermediate step before the final answer in the data, we prompt it with the problem only. See Figure 5c for prompting configuration in this case. We use final answer prediction correctness for this experiment to measure model performance. As displayed in Figure 2(d)2(e)2(f), in this setting, for all three base models, models that are fine-tuned/trained from scratch on SER got 0 final answers correct consistently. The results of SER show that LLMs don’t look forward during training, i.e. fail to learn to use intermediate supervision on the right in training.
@@
We also note that when both the problem and all intermediate proof steps (except for the final answer) are provided as prompts to the models, SEQ models still outperform PSER models in terms of final answer correctness, as shown in Figure 2(d)2(e)2(f). See Figure 5d, 5e for prompting configuration. By giving the correct intermediate steps as prompts to both SEQ and PSER models, we eliminate the snowballing effect during model generation.
@@
SER models learn spurious dependency.  In this experiment, we test our SEQ and SER models for the multiplication task. The testing prompt of the model is the problem only, and the models are expected to generate the full proof. See Figure 5a, 5c. We use the Step 1 correctness metric, i.e. checking the correctness of Step 1 in the models’ generated output. Step 1 is the decomposition step that splits the 4-digit number into simple terms, e.g. 2345 to 2*1000+3*100+4*10+5*1. During the training, the model should be able to learn this step quickly given the problem due to its simplicity. Indeed, from Figure 3(a)3(b), we see that for SEQ models based on pre-trained models, they learn Step 1 almost perfectly after the first 2500 steps. For the randomly initialized Gemma-2b trained from scratch on SEQ, the model performance measured by the Step 1 correctness metric quickly reaches more than 90% and is approaching perfect scores followed by some fluctuations, as shown by Figure 3(c). For SER models, however, they converge much slower measured by the Step 1 correctness metric in the early training stage compared to SEQ models based on the evidences shown in Figure 3(a)3(b)3(c). This slow convergence in the early training stage shows that SER models try hard to fit the spurious dependency of Step 1 on all other proof steps at first. They need more training steps to realize that Step 1 does not depend on all the other proof steps, including the final answer.
@@ 5 Other Results and Discussion
Bias in pre-training data is not the major cause of the order effect.  One might argue that the order bias in the pre-training dataset could potentially be the primary cause of the order effects observed in this study. However, for the multiplication task, we conduct controlled experiments on both pre-trained models and randomly initialized model, thereby ruling out this possibility given the results as shown in Figure 2(f).
SEQ is better than PSER even when problem and proof steps are masked during training.  In this experiment, we don’t use our 9 models mentioned before. We fine-tune pre-trained Gemma-2b on SEQ and PSER but with all tokens except final answers masked out. Therefore the only input-output pair that enters into the loss function is (problem+intermediate proof steps, final answer). During testing, we prompt both models with the problem and the intermediate proof steps. See Figure 5d,5e for the prompting configuration. Figure 10(b) shows that the model fine-tuned on masked SEQ outperforms the model fine-tuned scratch on masked PSER measured by final answer correctness for the pre-trained case.
@@
Discussion: artificiality of the dataset. It is important to note that the datasets used in this study for training and evaluating models are artificially constructed. Ideally, we would prefer to train and evaluate models on graduate- and research-level mathematical theorems and proofs. However, although state-of-the-art (SOTA) models can partially reorder proofs into a more intuitively sequential structure, there remains no reliable method for verifying the correctness of model-generated proofs. To illustrate the widespread nature of ordering issues in mathematical texts, we include in the appendix a list of proofs exhibiting the Dependency-Inversion Issue (DII), drawn from graduate-level mathematics textbooks. These constitute 17.3 percent of the theorems with nontrivial proofs in the first two chapters of the examined textbook, highlighting the severity of ordering problems in advanced mathematical reasoning.
@@ 6 Conclusion and Future work
In conclusion, our study highlights the significant impact of order for each data sample on LLMs’ training efficiency for the proof generation task. We claim that the optimal order of a data sample, which we call the intuitively sequential order, is to put intermediate supervision toward a particular proof step on the left for all proof steps in a data sample. We show the two causes for this order effect: given suboptimal data order, training inefficiency arises from (i) inability to look forward during training, i.e. learning to use future intermediate supervision to predict the current token, (ii) propensity to learn spurious token dependencies in the early training stage. Our experiments on the theorem-proving task and 4-by-4-digit multiplication provide strong support for these explanations, highlighting the importance of finding the optimal data order in model training. We explicitly define a particular kind of order issue, Delayed Insight Issue, that frequently appear in advanced math textbooks, and we present a list of proofs that suffer from DII in the appendix. Notably, 17.3 percent of theorems with non-trivial proofs in the first two chapters of the textbook we analyze suffer from DII.
A promising direction for future work is to investigate this claim during the pre-training stage of LLMs for reasoning tasks. Another future work is to employ machine learning methods to discover optimal data order. This holds significance on its own, independent of the gain it can bring to model training. For example, suppose the true supervision structures are discovered for mathematical proofs and those proofs are reordered accordingly. In that case, it might facilitate the learning process of students, making it linear and intuitive.
@@ Appendix A Appendix
A.1 Limitation
In this paper, we conduct experiments on models with 2B and 7B parameters. Due to computational constraints, we did not extend our experiments to larger models or incorporate additional ordering variations. As discussed in Section 5, we didn’t train and evaluate models on graduate or research-level math-proving datasets, since there is no direct way to check the correctness of the model-generated proofs.
A.2 Delayed Insight Issue: A common type of order issue that occurs frequently in graduate-level math textbooks
To illustrate DII, we compile a list of 9 cases from the first two chapters of a classic graduate-level textbook: Real Analysis, Modern Techniques and Their Applications authored by Gerald B. Folland (Folland1984RealAM). We find 17.3 percent of the theorems with non-trival proofs in the first two chapters of the textbook suffer from Delayed Insight Issue as defined in the main text.
A.2.1 Folland, Chapter 1, Proposition 1.3
Example 1 Proposition. If $A$ is countable, then $\bigotimes_{\alpha\in A}M_{\alpha}$ is the $\sigma$-algebra generated by $\{\prod_{\alpha\in A}E_{\alpha}:E_{\alpha}\in M_{\alpha}\}$ Proof: If $E_{\alpha}\in M_{\alpha}$, then $\pi_{\alpha}^{-1}(E_{\alpha})=\prod_{\beta\in A}E_{\beta}$ where $E_{\beta}=X_{\beta}$ for $\beta\neq\alpha$; on the other hand, $\prod_{\alpha\in A}E_{\alpha}=\cap_{\alpha\in A}\pi^{-1}_{\alpha}(E_{\alpha})$. The result therefore follows from Lemma 1.1.
Here, Lemma 1.1 serves as the major lemma, and the previous steps set up the stage for applying Proposition 1.4. Note that it is natural to first come up with the idea of applying Lemma 1.1 given the theorem.
A.2.2 Folland, Chapter 1, Proposition 1.5
Example 2 Proposition. Let $X_{1},..,X_{n}$ be metric spaces and let $X=\prod_{1}^{n}X_{j}$, equipped with the product metric. Then $\bigotimes_{1}^{n}B_{X_{j}}\subset B_{X}$. If the $X_{j}$’s are separable, then $\bigotimes_{1}^{n}B_{X_{j}}=B_{X}$ Proof: …… Moreover, the set of the set of points in $X$ whose jth coordinate is in $C_{j}$ for all j is a countable dense subset of $X$, and the balls of radius $r$ in $X$ are merely products of the balls of radius r in the $X_{j}$’s. It follows that $B_{X_{j}}$ is generated by $\mathbb{E}_{j}$ and $B_{X}$ is generated by $\prod^{n}_{1}E_{j}:E_{j}\in\mathbb{E}_{j}$. Therefore $B_{X}=\bigotimes_{1}^{n}B_{X_{j}}$ by Proposition 1.4
Here, Proposition 1.4 serves as the major lemma, and the previous steps set up the stage for applying Proposition 1.4. Note that it is natural to first come up with the idea of applying Proposition 1.4 given the theorem.
A.2.3 Folland, Chapter 1, Proposition 1.16
Example 3 Proposition. If $F:\mathbb{R}\mapsto\mathbb{R}$ is any increasing, right continuous function, there is a unique Borel measure $\mu_{F}$ on $\mathbb{R}$ such that $\mu_{F}((a,b])=F(b)-F(a)$ for all $a,b$. If $G$ is another such function, we have $\mu_{F}=\mu_{G}$ iff $F-G$ is a constant. Proof: Each $F$ induces a premeasure on $A$ by Proposition 1.15. It is clear that $F$ and $G$ induces the same premeasure iff $F-G$ is constant, and that these premeasures are $\sigma$-finite (since $\mathbb{R}=\cup^{\infty}_{-\infty}(j,j+1]$). The first two assertions therefore follow from Theorem 1.14.
Here, Theorem 1.14 serves as the major lemma, and the previous steps set up the stage for applying Theorem 1.14. Note that it is natural to first come up with the idea of applying Theorem 1.14 given the theorem.
A.2.4 Folland, Chapter 2, Proposition 2.6
Example 4 Proposition. If $f,g:X\mapsto\mathbb{C}$ are $M$-measurable, then so are $f+g$ and $fg$ Proof: Define $F:X\mapsto\mathbb{C}\times\mathbb{C}$, ϕ:×:italic-$\phi:\mathbb{C}\times\mathbb{C}\mapsto\mathbb{C}$ by $F(x)=(f(x),g(x))$, $\phi(z,w)=z+w$, $\psi(z,w)=zw$. Since $B_{\mathbb{C}\times\mathbb{C}}=B_{\mathbb{C}}\bigotimes B_{\mathbb{C}}$ by Proposition 1.5, $F$ is $(M,B_{\mathbb{C}\times\mathbb{C}})$-measurable by Proposition 2.4.
Here, Proposition 2.4 serves as the major lemma, and the previous steps set up the stage for applying Theorem 1.14. Note that it is natural to first come up with the idea of applying Proposition 2.4 given the theorem, and then make the constructions as in the early part of the proof.
A.2.5 Folland, Chapter 2, Corollary 2.17
Example 5 Corollary. If $\{f_{n}\}\subset L^{+}$, $f\in L^{+}$, and $f_{n}(x)$ increases to $f(x)$ for a.e $x$, then $\int f=\lim\int f_{n}$ Proof: If $f_{n}(x)$ increases to $f(x)$ for $x\in E$ where $\mu(E^{c})=0$, then $f-f\chi_{E}=0$ a.e. and $f_{n}-f_{n}\chi_{E}=0$ a.e., so by the monotone convergence theorem, $\int f=\int f\chi_{E}=\lim\int f_{n}\chi_{E}=\lim\int f_{n}$
For this problem, the first instinct after looking at the corollary should be to apply the monotone convergence theorem. The earlier steps explicitly state how the monotone convergence theorem should be applied to the given corollary.
A.2.6 Folland, Chapter 2, Fatou’s Lemma 2.18
Example 6 Fatou’s Lemma. If $\{f_{n}\}$ is any sequence in $L^{+}$, then $\int(\lim\inf f_{n})\leq\lim\inf\int f_{n}$ Proof: For each $k\geq 1$ we have $\inf_{n\geq k}f_{n}\leq f_{j}$ for $j\geq k$, hence $\int\inf_{n\geq k}f_{n}\leq\int f_{j}$ for $j\geq k$, hence $\int\inf_{n\geq k}f_{n}\leq\inf_{j\geq k}\int f_{j}$. Now let $k\mapsto\infty$ and apply monotone convergence theorem: $\int(\lim\inf f_{n})=\lim_{k\mapsto\infty}\int(\inf_{n\geq k}f_{n})\leq\lim\inf\int f_{n}$
For this problem, the first instinct after looking at the lemma should be to apply the monotone convergence theorem. The earlier steps explicitly state how to transform the problem to a state so that one can apply the monotone convergence theorem.
A.2.7 Folland, Chapter 2, Theorem 2.25
Example 7 Theorem. Suppose that $\{f_{j}\}$ is a sequence in $L^{1}$ such that $\Sigma_{1}^{\infty}\int|f_{j}|<\infty$. Then $\Sigma_{1}^{\infty}f_{j}$ converges a.e. to a function in $L^{1}$, and $\int\Sigma_{1}^{\infty}f_{j}=\Sigma_{1}^{\infty}\int f_{j}$ Proof: By Theorem 2.15, $\int\Sigma_{1}^{\infty}|f_{j}|=\Sigma_{1}^{\infty}\int|f_{j}|<\infty$, so that function $g=\Sigma_{1}^{\infty}|f_{j}|$ is in $L^{1}$. In particular, by Proposition 2.20, $\Sigma_{1}^{\infty}|f_{j}(x)|$ is finite for a.e. $x$, and for each such $x$ the series $\Sigma_{1}^{\infty}f_{j}(x)$ converges. Moreover, $|\Sigma_{1}^{n}f_{j}|\leq g$ for all $n$, so we can apply the dominated convergence theorem to the sequence of partial sums to obtain $\int\Sigma_{1}^{\infty}f_{j}=\Sigma_{1}^{\infty}\int f_{j}$
For this problem, the first instinct after looking at the lemma should be to apply the dominated convergence theorem. The earlier steps explicitly state how to make sure the conditions of the dominated convergence theorem are satisfied by the statement.
A.2.8 Folland, Chapter 2, Theorem 2.27
[…]
For this problem, the first instinct after looking at the lemma should be to apply the dominated convergence theorem. The earlier steps explicitly state how to make sure the conditions of the dominated convergence theorem are satisfied by the statement.
A.2.9 Folland, Chapter 2, Theorem 2.47
[…]
For this problem, without first having a clear goal as expressed in the previous equation, one will be completely confused for why we need to do all the estimation at the beginning.
A.3 Prompting configuration for testing
See Figure 5.
A.4 Data Examples in training